# GASP: Gradient-Aware Shortest Path Algorithm for Boundary-Confined 2-Manifold Reeb Graph Visualization

### v1

@@ 7 Conclusion and Future Work
Efficient calculation of the Reeb graphs has been a well-studied problem. However, the actual visualization of Reeb graphs is understudied. When abstracting and representing data, accurately representing it is of the utmost importance [3]. The geometric barycenter algorithm used in the state-of-the-practice tool TTK, has some drawbacks, including producing arcs outside of the model, which are longer than necessary and do not optimally follow the function gradient. GASP demonstrates significant qualitative and quantitative improvements over the geometric barycenter algorithm by constraining arcs to the model’s surface and utilizing a gradient-aware shortest paths.
Despite the improvement provided by GASP, the approach still presents some limitations. First, the interior method assumes that the contour is approximately planar. As the shape of the contour becomes less planar, it increases the likelihood that the interior approach will produce a low-quality Reeb graph arc. In addition, we have implemented GASP using Python. In the future, we will investigate the extension of our methods to 3D meshes, with challenges associated with Reeb graph representation arising from tetrahedra and isosurface evolution.

### v2

@@ 7 Conclusion and Future Work
Efficient calculation of the Reeb graphs has been a well-studied problem. However, the actual visualization of Reeb graphs is understudied. When abstracting and representing data, accurately representing it is of the utmost importance [3]. The geometric barycenter algorithm used in the state-of-the-practice tool TTK has some drawbacks, including producing arcs outside of the model that are longer than necessary and do not optimally follow the function gradient. GASP demonstrates significant qualitative and quantitative improvements over the geometric barycenter algorithm by constraining arcs to the model’s surface and utilizing a gradient-aware shortest paths.
Despite the improvement provided by GASP, the approach still presents some limitations. First, the interior method assumes that the contour is approximately planar and the gradient of the function is well defined. For the height function, the contours are planar and the gradient is well-defined. However, for non-planar contours, determining interior points in the embedding space is not straightforward and the gradient is ill-defined. As the contour becomes less planar, the likelihood increases that the interior approach will produce a low-quality Reeb graph arc due to inaccurate approximation of interior points. Second, although GASP can technically work on any Morse function, we evaluated it only for the height and geodesic functions. Additional examples using a function defined as the sum of 3D Gaussians centered at random vertices are provided in the appendix, where our method produces Reeb graphs of similar or better quality compared to TTK. In the future, we will investigate the extension of GASP methods to 3D meshes, with challenges associated with Reeb graph representation arising from tetrahedra and isosurface evolution.
Source code and data for the project are available at https://github.com/shape-vis/gasp.
Acknowledgments
This project is supported in part by U.S. Department of Energy (DOE) RAPIDS-2 SciDAC project under contract number DE-AC0500OR22725 and the National Science Foundation (IIS-2316496).